# How to Understand “Support”? An Implicit-enhanced Causal Inference Approach for Weakly-supervised Phrase Grounding

### 29 Feb 2024 v1

@@ Abstract
Jiamin Luo, Jianing $*{}^{*}$thanks: $*{}^{*}$ Equal contribution, Jingjing ${}^{\dagger}$thanks: ${}^{\dagger}$ Corresponding author, Guodong Zhou
@@ 1.   Introduction
Despite this, these studies fail to deeply explore the semantic nature of phrases that some phrases often exhibit the implicit and intricate semantics, rendering it arduous for models to establish correct connections with image regions. Take Figure 1 as an example, the phrase “support” necessitates the integration of the commonsense knowledge to precisely find its corresponding image region, i.e., red boxes. In this study, we refer to such correspondence as one type of implicit phrase-region matching relations [^1] In this paper, this commonsense-involved implicit relation is named as “commonsense understanding”. Besides, we also propose another three implicit relations as shown in Figure 3., and the grasp of such implicit relations serves as a valuable evaluation of the model capability in understanding the deep multimodal semantics. Additionally, the weakly-supervised setting of WPG could bring the supervised noise problem Xiao et al. (2017) which may lead to more difficulties in capturing such implicit information, making it urgent to address the implicit relations problem in WPG. In this study, we contend that capturing such relations at least faces two main challenges.
@@ 4.2.   Baselines
We choose several state-of-the-art baselines for WPG to compare performance with our IECI approach, described as follows. KAC-Net Chen et al. (2018) explores the consistency in visual and language as complementary external knowledge. ARN Liu et al. (2019a) builds the correspondence between image region and query in an adaptive manner. KPRN Liu et al. (2019b) models the relation between target and contextual entities. InfoGround Gupta et al. (2020) leverages contrastive learning to maximize a lower bound between region features and contextualized word representations. ALBEF Li et al. (2021) introduces a contrastive loss to align the image and text representations before fusing them through cross-modal attention. CL&KD Wang et al. (2021a) focuses on distilling knowledge from a generic object detector under the framework of contrastive learning. ReIR Liu et al. (2021) incorporates coarse-to-fine object refinement and entity relation modeling into a two-stage deep network. Moreover, an advanced multimodal pre-training model BLIP Li et al. (2022) is leveraged to compare with IECI. Here, it should be noted that BLIP can not directly fine-tune the WPG task, since BLIP is pre-trained with the coarse-grained sentence-image pairs. Thus, we treat the phrase-region pairs as the coarse-grained pairs to fine-tune the WPG task.
@@ 4.3.   Implementation Details and Metrics
In our experiments, we leverage open-source codes to obtain experimental results of all the baselines on both Implicit and Explicit datasets, and re-implement the results of several baselines by their open-source codes on Full datasets. The hyper-parameters of these baselines reported by their public papers maintain the same setting, and the others are tuned according to the validation set. Specifically, for phrase encoder, the parameters of BERT are following Devlin et al. (2019). For region encoder, we utilize a pre-trained Faster R-CNN on the Visual Genome Krishna et al. (2017) dataset to generate a maximum number of 100 object regions. For IECI, we employ the Adam optimizer with an initial learning rate of 1e-5 for training. The regularization weight of parameters is 1e-4. The hyper-parameter $\alpha$ is set to be 0.1 and the batch size is set to be 64. The layers of the self-sampling attention and cross-sampling attention in IDA are both set to be 6. The overall training parameters of our approach are 0.13B. To facilitate the corresponding research in this direction, all codes alongside our implicit-enhanced dataset will be released via Github.
Besides, the WPG performance is evaluated using Recall@$k$, $k\in(1,5)$, which measures the fraction of phrases for which the ground-truth bounding box exhibits an IoU $\geqslant$ 0.5 with any of the top-$k$ predicted boxes. Moreover, $t$-test [^2] https://www.scipy.org/ is used to evaluate the significance of the performance difference.
@@ 5.1.   Experimental Results
Table 1 shows the performance comparison of different approaches. From this table, we can see that: 1) The performances of all approaches on the Implicit dataset are consistently lower than the Explicit dataset. This indicates that modeling the implicit relations is more challenging than the explicit, and encourages us to consider such implicit relations problem in WPG. 2) The performance improvements of our IECI approach on the Implicit dataset are larger than those on the Explicit dataset. This indicates that IECI can effectively address the challenge of highlighting the implicit relations beyond the explicit. Furthermore, our IECI approach consistently performs better than all other approaches. Impressively, compared to the best-performing ReIR approach, our IECI approach achieves the average R@1 improvements of 2.83%, 2.29% and 2.34% on all three Implicit, Explicit and Full datasets, respectively. Significance test shows that all these improvements are significant ($p$-value < 0.01). These demonstrate that IECI can better highlight the implicit relations and meantime maintain the performance of the explicit. Particularly, we observe that the performance of BLIP is quite worse than IECI. This is reasonable since BLIP is pre-trained on the coarse-grained sentence-images pairs, thus is not suitable for aligning the fine-grained phrase-region pairs. 3) Our IECI approach trained on both the Flickr30K and COCO datasets still outperforms all other approaches. This justifies the robustness of IECI, and again encourages us to consider the important implicit relations in the WPG task.
@@ 5.3.   Comparison with Multimodal LLMs
Recently, the emergence of large language models (LLMs) has demonstrated their remarkable abilities across various fields and tasks. Since GPT-4 is closed-sourced and currently cannot evaluate the multimodal tasks, we choose two advanced open-sourced multimodal LLMs, i.e., MiniGPT4-13B Zhu et al. (2023) and LLaVA-13B Liu et al. (2023), to verify the effectiveness of IECI. Specifically, we first randomly select 30 sentence-image examples with the implicit and explicit relations for comparisons on the WPG task. Then, we select two well-studied evaluation methods for LLMs as follows: 1) Zero-Shot (ZS). Following the ZS setting for evaluating LLMs proposed by Bang et al. (2023), given a prompt including the task instructions, sentence-image pair and the bounding box list of each phrase, we utilize multimodal LLMs to “generate the bounding boxes from the list that can reflect the given phrase”. 2) In-Context Learning (ICL). Following the ICL setting for evaluating LLMs proposed by Li et al. (2023b), building upon the ZS, we introduce some additional demonstration examples “phrase, boxes;…;phrase, boxes” as the prompt. As shown in Figure 4, we observe that IECI outperforms the multimodal LLMs by a large margin, indicating that the multimodal LLMs still face challenges in understanding the deep multimodal semantics. Moreover, we find that the performance of ZS is better than ICL, since the image encoders BLIP-2 and CLIP within MiniGPT4-13B and LLaVA-13B lack the ability of ICL as reported in Li et al. (2023a) and Radford et al. (2021).

### 04 Mar 2024 v2

@@ Abstract
Jiamin $*{}^{*}$, Jianing $*{}^{*}$thanks: $*{}^{*}$qual Contribution., Jingjing ${}^{\dagger}$thanks: ${}^{\dagger}$orresponding Author: Jingjing Wang., Guodong Zhou
@@ 1.   Introduction
Despite this, these studies fail to deeply explore the semantic nature of phrases that some phrases often exhibit the implicit and intricate semantics, rendering it arduous for models to establish correct connections with image regions. Take Figure 1 as an example, the phrase “support” necessitates the integration of the commonsense knowledge to precisely find its corresponding image region, i.e., red boxes. In this study, we refer to such correspondence as one type of implicit phrase-region matching relations [^1] In this paper, this commonsense-involved implicit relation is named as “commonsense understanding”. Besides, we also propose another three implicit relations as shown in Figure 3., which could be defined as phrases in the sentence that are not specific and explicit nominal phrases of the objects in the image, and the grasp of such implicit relations serves as a valuable evaluation of the model capability in understanding the deep multimodal semantics. Additionally, the weakly-supervised setting of WPG could bring the supervised noise problem Xiao et al. (2017) which may lead to more difficulties in capturing such implicit information, making it urgent to address the implicit relations problem in WPG. In this study, we contend that capturing such relations at least faces two main challenges.
@@ 4.2.   Baselines
We choose several state-of-the-art baselines for WPG to compare performance with our IECI approach, described as follows.
  • KAC-Net Chen et al. (2018) explores the consistency in visual and language as complementary external knowledge.
  • ARN Liu et al. (2019a) builds the correspondence between image region and query in an adaptive manner.
  • KPRN Liu et al. (2019b) models the relation between target and contextual entities.
  • InfoGround Gupta et al. (2020) leverages contrastive learning to maximize a lower bound between region features and contextualized word representations.
  • ALBEF Li et al. (2021) introduces a contrastive loss to align the image and text representations before fusing them into cross-modal attention.
  • CL&KD Wang et al. (2021a) focuses on distilling knowledge from a generic object detector under the framework of contrastive learning.
  • ReIR Liu et al. (2021) incorporates coarse-to-fine object refinement and entity relation modeling into a two-stage deep network.
Moreover, an advanced multimodal pre-training model BLIP Li et al. (2022) is leveraged to compare with IECI. Here, it should be noted that BLIP can not directly fine-tune the WPG task, since BLIP is pre-trained with the coarse-grained sentence-image pairs. Thus, we treat the phrase-region pairs as the coarse-grained pairs to fine-tune the WPG task.
@@ 4.3.   Implementation Details and Metrics
In our experiments, we leverage open-source codes to obtain experimental results of all the baselines on both Implicit and Explicit datasets, and re-implement the results of several baselines by their open-source codes on Full datasets. The hyper-parameters of these baselines reported by their public papers maintain the same setting, and the others are tuned according to the validation set. Specifically, for phrase encoder, the parameters of BERT are following Devlin et al. (2019). For region encoder, we utilize a pre-trained Faster R-CNN on the Visual Genome Krishna et al. (2017) dataset to generate a maximum number of 100 object regions. For IECI, we employ the Adam optimizer with an initial learning rate of 1e-5 for training. The regularization weight of parameters is 1e-4. The hyper-parameter $\alpha$ is set to be 0.1 and the batch size is set to be 64. The layers of the self-sampling attention and cross-sampling attention in IDA are both set to be 6. The overall training parameters of our approach are 0.16B. All the baselines are reproduced using PyTorch on a machine equipped with an NVIDIA GeForce RTX 3090, an Intel(R) Xeon(R) E5-2650 v4 CPU (2.20 GHz), CUDA version 11.7, and the PyTorch 1.7.1 library with Python 3.6.13, running on Ubuntu 20.04.1 LTS. To facilitate the corresponding research in this direction, all codes alongside our implicit-enhanced dataset will be released via Github  [^2] Github link will come soon..
Besides, the WPG performance is evaluated using Recall@$k$, $k\in(1,5)$, which measures the fraction of phrases for which the ground-truth bounding box exhibits an IoU $\geqslant$ 0.5 with any of the top-$k$ predicted boxes. Moreover, $t$-test [^3] https://docs.scipy.org/doc/scipy/reference/stats.html is used to evaluate the significance of the performance difference.
@@ 5.1.   Experimental Results
Table 1 shows the performance comparison of different approaches. From this table, we can see that: 1) The performances of all approaches on the Implicit dataset are consistently lower than the Explicit dataset. This indicates that modeling the implicit relations is more challenging than the explicit, and encourages us to consider such implicit relations problem in WPG. Besides, we also observe that the performance of both Implicit and Explicit datasets is improved, and we analyze the reason is that there may be some overlap between the implicit and explicit relations in the current dataset, leading to performance improvements in explicit relations when employing causal inference. 2) The performance improvements of our IECI approach on the Implicit dataset are larger than those on the Explicit dataset. This indicates that IECI can effectively address the challenge of highlighting the implicit relations beyond the explicit. Furthermore, our IECI approach consistently performs better than all other approaches. Impressively, compared to the best-performing ReIR approach, our IECI approach achieves the average R@1 improvements of 2.83%, 2.29% and 2.34% on all three Implicit, Explicit and Full datasets, respectively. Significance test shows that all these improvements are significant ($p$-value < 0.01). These demonstrate that IECI can better highlight the implicit relations and meantime maintain the performance of the explicit. Particularly, we observe that the performance of BLIP is quite worse than IECI. This is reasonable since BLIP is pre-trained on the coarse-grained sentence-images pairs, thus is not suitable for aligning the fine-grained phrase-region pairs. 3) Our IECI approach trained on both the Flickr30K and COCO datasets still outperforms all other approaches. This justifies the robustness of IECI, and again encourages us to consider the important implicit relations in the WPG task.
((a))
@@ 5.3.   Comparison with Multimodal LLMs
Recently, the emergence of large language models (LLMs) has demonstrated their remarkable abilities across various fields and tasks. Since GPT-4 is closed-sourced and currently cannot evaluate the multimodal tasks, we choose two advanced open-sourced multimodal LLMs, i.e., MiniGPT4-13B Zhu et al. (2023) and LLaVA-13B Liu et al. (2023), to verify the effectiveness of IECI. Specifically, we first randomly select 30 sentence-image examples with the implicit and explicit relations for comparisons on the WPG task. Then, we select two well-studied evaluation methods for LLMs as follows: 1) Zero-Shot (ZS). Following the ZS setting for evaluating LLMs proposed by Bang et al. (2023), given a prompt including the task instructions, sentence-image pair and the bounding box list of each phrase, we utilize multimodal LLMs to “generate the bounding boxes from the list that can reflect the given phrase”. 2) In-Context Learning (ICL). Following the ICL setting for evaluating LLMs proposed by Li et al. (2023b), building upon the ZS, we introduce one additional demonstration example “phrase, boxes;…;phrase, boxes” as the prompt. As shown in Figure 4, we observe that IECI outperforms the multimodal LLMs by a large margin, indicating that the multimodal LLMs still face challenges in understanding the deep multimodal semantics. Moreover, we find that the performance of ZS is better than ICL, since the image encoders BLIP-2 and CLIP within MiniGPT4-13B and LLaVA-13B lack the ability of ICL as reported in Li et al. (2023a) and Radford et al. (2021).